# NoisyRollout: Reinforcing Visual Reasoning with Data Augmentation

### v3

@@ Appendix B Unsuccessful Attempts
Reward penalty on noisy subgroup. We experimented with applying explicit reward penalties (e.g., $-0.1$) to all noisy rollouts, aiming to encourage the model to better capture contrastive learning signals. However, this approach quickly led to training divergence. Rather than improving its core reasoning and perception abilities, the policy model learned to distinguish between clean and noisy rollouts. As a result, the noisy rollouts rapidly became highly “off-policy”, since the model could easily identify. This distributional mismatch destabilized training and undermined the learning.
Appendix C Training Procedure
To better illustrate the workflow of our method, a simplified overview is provided in Algorithm 1.
Algorithm 1 NoisyRollout: Noisy Reinforcement Fine-Tuning
2:for $t=1$ to $t_{\text{max}}$ do
3:     Sample batch $(I,\mathbf{q})\sim p_{\mathcal{D}}$
4:     Set noise strength $\alpha_{t}=\eta(\alpha_{0},t,t_{\text{max}})$triangleright Annealing schedule
5:     Generate distorted images $\tilde{I}=T_{\alpha_{t}}(I)$
6:     Sample $\{\mathbf{o}_{j}\}_{j=1}^{n_{1}}$ from $\pi_{\theta_{\text{old}}}(\mathbf{o}\mid I,\mathbf{q})$triangleright Clean rollouts
7:     Sample $\{\mathbf{o}_{k}\}_{k=n_{1}+1}^{n_{1}+n_{2}}$ from $\pi_{\theta_{\text{old}}}(\mathbf{o}\mid\tilde{I},\mathbf{q})$triangleright Noisy rollouts
10:     Update policy using:
12:     $\theta\leftarrow\theta-\nabla_{\theta}\mathcal{J}(\theta)$triangleright Update conditioned on clean images only
13:     $\theta_{\text{old}}\leftarrow\theta$triangleright Update old policy parameters
14:end for
Appendix D Evaluating the Perception Quality of Reasoning Traces
@@
Visual Information Extraction Prompt Extract all visual perception and information recognition components from the following reasoning trace. Original question: {question} Reasoning trace: {reasoning} Your task is to extract and summarize ONLY the parts that relate to visual perception, information extraction, and understanding of visual elements from the image. This includes: 1. Any measurements, dimensions, or numerical values extracted from the image 2. Description of visual elements like shapes, objects, positions, or spatial relationships 3. Recognition of text, symbols, diagrams, or graphs from the image 4. Any visual features mentioned or used in the reasoning Format your response with the tag: <<<visual_perception>>> [Extracted visual information here] <<</visual_perception>>> Include ONLY visual perception elements, not mathematical reasoning that happens after the information is extracted. If there are no clear visual perception elements, respond with “No clear visual perception elements identified.” Visual Perception Comparison Prompt Compare the quality of visual perception between two models based on the image and the original question. Original question: {question} Visual perception from Model A: {visual A} Visual perception from Model B: {visual B} Your task is to determine which model better captures and correctly extracts visual information from the image. Compare their visual perception quality based on: Accuracy of visual information extraction (measurements, shapes, relationships) Complete identification of all relevant visual elements Proper recognition of visual information required to solve the problem Score both models and determine the winner: If Model A demonstrates significantly better perception than Model B, respond: <<<result>>>A<<</result>>>; if Model B demonstrates significantly better perception than Model A, respond: <<<result>>>B<<</result>>>; if both models show similar quality of visual perception, respond: <<<result>>>tie<<</result>>>. Now: 1. identify what visual information is required to solve this problem. 2. analyze how each model perceives this information. 3. provide your comparative judgment with specific reasons. 4. provide your <<<result>>> tag with exactly A, B, or tie.
@@
Appendix E Detailed Related Work
@@
In the multimodal domain, research on leveraging RL to enhance VLMs’ reasoning capabilities remains in early stages. Some approaches explore using generative reward models [84, 66] to enhance VLMs’ general capability, but these typically require powerful closed-source models for training data generation. Recent work including LMM-R1 [51], Vision-R1 [21], R1-V [7] and OpenVLThinker [13] has applied R1-type RL to VLMs in diverse specific subdomains like geometry problems and object counting tasks [50, 10, 12, 33, 63, 24]. Further more, pilot studies [48, 46] further extend large-scale rule-based RL to broader multimodal mathematical reasoning, demonstrating significant performance gains without relying on in-domain training data.
Appendix F Broader Impacts
@@
Appendix G Limitations
Despite easy-to-adopt designs and promising empirical results, our study has several limitations. First, due to computational constraints, our experiments are limited in scale: we primarily explore model sizes up to $32$B parameters and training dataset scales in the order of a few thousand samples. Future work should validate and extend our findings using significantly larger-scale training scenarios—such as models with $72$B parameters or training datasets in the range of hundreds of thousands of sample. Second, NoisyRollout is applied during the RL fine-tuning phase of an already pre-trained VLM. A more fundamental, but vastly more complex, direction would be to explore how the principles of NoisyRollout (i.e., learning from noisy signals in RL) could be integrated into the large-scale pre-training phase of the VLM itself. Finally, while empirically effective, our study lacks a formal theoretical analysis of how NoisyRollout, with its specific hybrid trajectory mixing and noise annealing, affects the exploration-exploitation trade-off and the convergence properties of the RL algorithm. It’s unclear if the introduced noise guarantees broader state-space coverage in a principled way or if certain noise characteristics could inadvertently hinder convergence.
Appendix H Case Study
@@ Appendix I Templates
Appendix J Detailed Methodology for Gradient Contribution Analysis
@@
Appendix K Supplementary Implementation Details
@@
Sigmoid Steepness ($\lambda$) $t_{\text{max}}/2\ (30)$
@@
Sigmoid Steepness ($\lambda$) $t_{\text{max}}/2\ (60)$
@@
Training Episodes $7$
Total Optimization Steps ($t_{\text{max}}$) $84$

### v4

@@ Appendix B Unsuccessful Attempts
Reward penalty on noisy subgroup. We experimented with applying explicit reward penalties (e.g., 0.1-0.1) to all noisy rollouts, aiming to encourage the model to better capture contrastive learning signals. However, this approach quickly led to training divergence. Rather than improving its core reasoning and perception abilities, the policy model learned to distinguish between clean and noisy rollouts. As a result, the noisy rollouts rapidly became highly “off-policy”, since the model could easily identify. This distributional mismatch destabilized training and undermined the learning.
Appendix C Evaluating the Perception Quality of Reasoning Traces
Visual Information Extraction Prompt Extract all visual perception and information recognition components from the following reasoning trace. Original question: {question} Reasoning trace: {reasoning}
@@ Appendix C Evaluating the Perception Quality of Reasoning Traces
Your task is to extract and summarize ONLY the parts that relate to visual perception, information extraction, and understanding of visual elements from the image. This includes: 1. Any measurements, dimensions, or numerical values extracted from the image 2. Description of visual elements like shapes, objects, positions, or spatial relationships 3. Recognition of text, symbols, diagrams, or graphs from the image 4. Any visual features mentioned or used in the reasoning Format your response with the tag: <<visual_perception>> [Extracted visual information here] <</visual_perception>> Include ONLY visual perception elements, not mathematical reasoning that happens after the information is extracted. If there are no clear visual perception elements, respond with “No clear visual perception elements identified.” Visual Perception Comparison Prompt Compare the quality of visual perception between two models based on the image and the original question. Original question: {question} Visual perception from Model A: {visual A}
Visual perception from Model B: {visual B}
Your task is to determine which model better captures and correctly extracts visual information from the image. Compare their visual perception quality based on: Accuracy of visual information extraction (measurements, shapes, relationships) Complete identification of all relevant visual elements Proper recognition of visual information required to solve the problem Score both models and determine the winner: If Model A demonstrates significantly better perception than Model B, respond: <<result>>A<</result>>; if Model B demonstrates significantly better perception than Model A, respond: <<result>>B<</result>>; if both models show similar quality of visual perception, respond: <<result>>tie<</result>>. Now: 1. identify what visual information is required to solve this problem. 2. analyze how each model perceives this information. 3. provide your comparative judgment with specific reasons. 4. provide your <<result>> tag with exactly A, B, or tie.
@@
Appendix D Detailed Related Work
@@
In the multimodal domain, research on leveraging RL to enhance VLMs’ reasoning capabilities remains in early stages. Some approaches explore using generative reward models [95, 76] to enhance VLMs’ general capability, but these typically require powerful closed-source models for training data generation. Recent work including LMM-R1 [56], Vision-R1 [23], R1-V [7] and OpenVLThinker [13] has applied R1-type RL to VLMs in diverse specific subdomains like geometry problems and object counting tasks [55, 10, 12, 39, 37, 73, 26]. Further more, pilot studies [53, 51] further extend large-scale rule-based RL to broader multimodal mathematical reasoning, demonstrating significant performance gains without relying on in-domain training data.
Data Augmentation in Visual Reinforcement Learning. Data augmentation has played a central role in visual reinforcement learning (RL) from pixels by improving sample efficiency and generalization without architectural changes. RAD (Reinforcement Learning with Augmented Data) first demonstrated that simple image transformations—random crop, translation, color jitter, and cutout—can substantially enhance policy learning, particularly in DMControl and ProcGen benchmarks [33]. CURL combined off-policy RL with contrastive learning on augmented views, improving representation quality and sample efficiency [68]. DrQ regularized Q-value targets and policy updates via random shifts, achieving stable gains with minimal computational overhead [30]; DrQ-v2 further refined training schedules, exploration, and multi-step targets to master high-dimensional control from pixels [85].
Beyond these baselines, several extensions explored stronger invariance and generalization. DrAC introduced actor–critic regularizers that promote consistency under augmentation and automated augmentation search for diverse environments [60]. SVEA mitigated instability under heavy augmentations by mixing augmented and clean views to reduce target variance [22]. Self-Predictive Representations (SPR) encouraged temporal consistency between augmented latent states to improve representation quality and low-data efficiency [63]. In the sim-to-real context, domain randomization perturbs textures, lighting, and geometry to improve transferability from simulation to real-world environments [71, 61].
While these studies have established augmentation as a key ingredient for robust visual control, most methods optimize control policies rather than multimodal reasoning. Current augmentation pipelines are designed to stabilize perception and dynamics modeling but rarely connect augmented visual representations to higher-level symbolic or linguistic reasoning. Bridging this gap will likely require augmentations that preserve both control-relevant and semantic information, enabling joint optimization for perception, reasoning, and alignment in future multimodal agents.
Appendix E Broader Impacts
@@
Appendix F Limitations
Despite easy-to-adopt designs and promising empirical results, our study has several limitations. First, due to computational constraints, our experiments are limited in scale: we primarily explore model sizes up to 3232B parameters and training dataset scales in the order of a few thousand samples. Future work should validate and extend our findings using significantly larger-scale training scenarios—such as models with 7272B parameters or training datasets in the range of hundreds of thousands of sample. Second, NoisyRollout is applied during the RL fine-tuning phase of an already pre-trained VLM. A more fundamental, but vastly more complex, direction would be to explore how the principles of NoisyRollout (i.e., learning from noisy signals in RL) could be integrated into the large-scale pre-training phase of the VLM itself. Finally, while empirically effective, our study lacks a formal theoretical analysis of how NoisyRollout, with its specific hybrid trajectory mixing and noise annealing, affects the exploration-exploitation trade-off and the convergence properties of the RL algorithm. It’s unclear if the introduced noise guarantees broader state-space coverage in a principled way or if certain noise characteristics could inadvertently hinder convergence.
Appendix G Case Study
@@ Appendix H Templates
Appendix I Detailed Methodology for Gradient Contribution Analysis
@@
Appendix J Supplementary Implementation Details
@@
Sigmoid Steepness ($\lambda$) $30$
@@
Sigmoid Steepness ($\lambda$) $60$
@@
Training Episodes 77
Total Optimization Steps ($t_{\text{max}}$) $70$
@@
Appendix K Licenses
We use standard licenses from the community. We include the following licenses for the codes, datasets and models we used in this paper.
Datasets & Benchmarks:
Codes:
Models: